OpenStudy (anonymous):

The volumes of two similar solids is 32 cm3 and 864 cm3. What is the ratio of the corresponding sides

OpenStudy (anonymous):

ratio of their sides =cube root of (32/864)

OpenStudy (anonymous):

1/3

OpenStudy (shamim):

i think both solids r cube

OpenStudy (shamim):

is not it

OpenStudy (anonymous):

Apply similarity

OpenStudy (anonymous):

(volume factor) = (Scale factor)^3 And take cube root on both sides of the equation

OpenStudy (anonymous):

refer to the top / first 2 answers above

OpenStudy (anonymous):

@matricked , can you show me work?

ganeshie8 (ganeshie8):

if the side lengths are in ratio \(\large \mathbb{\frac{p}{q}}\), then the area will be in ratio \(\large \mathbb{\frac{p^2}{q^2}}\), and the volumes will be in ratio \(\large \mathbb{\frac{p^3}{q^3}}\),

ganeshie8 (ganeshie8):

working it backwards gives u the side lengths

ganeshie8 (ganeshie8):

\[\dfrac{p}{q} = \left(\dfrac{p^3}{q^3}\right)^{\dfrac{1}{3}}\]

OpenStudy (anonymous):

Thank you as always @ganeshie8 And everyone else Thank you soo much!!!!